OpenStudy (anonymous):

QUADRATIC EQUATION. Solve 5x2 = -30x - 65.

OpenStudy (anonymous):

I don't need solved all the way down to the roots. Just as simplified as possible.

OpenStudy (anonymous):

If you mean 5x2=5x^2 Then you can write it as 5x^2+30x+65=0 <=> x^2+6x+13=0 This has no solutions in real numbers as it's Delta which is equal to 6^2-4*13=-16 is negative. Thus there are no real solutions.

OpenStudy (anonymous):

No, I need the numbers filed in the quadratic equation. I have it down to this\[\frac{ -30 \pm \sqrt{900 - 1300} }{ 10 }\]

OpenStudy (anonymous):

You have a square root of a negative numbers. There are no real numbers that are square roots of negative values. If you do insisnt od doing it and you have it down to this just divide it by ten ( by a hundred in a square root)

OpenStudy (anonymous):

Well you have to use imaginary numbers..

OpenStudy (anonymous):

Then i am out of my league. Sorry for not being able to help.